OpenStudy (anonymous):

A large bookstore took in $3,000,000 in sales last year The ratio of book sales to other sales was 3 to 2. How much money in book sales did the store take in?

OpenStudy (anonymous):

3x + 2x = 3000000 x = 600000 so book sales 3 * 600000 = $1800000

OpenStudy (anonymous):

Just calculate the ratio. $3,000,000*3/5=$1,800,000